# Ali Dasdan KD Consulting Saratoga, CA, USA

### v2

@@ 5.5 What Would Work
Break Condition 3 (Dominance): Reduce existential stakes making defection dominant. Mechanisms might include: (1) credible international agreements reducing security competition, (2) economic restructuring reducing winner-take-all dynamics, or (3) demonstrated proof that advanced AI poses catastrophic risks even to developers. This is theoretically possible but requires fundamental shifts in international relations and economic structure.
Break Condition 4 (Pareto-Inefficiency): Accept the racing equilibrium rather than seeking coordination. Some argue racing toward advanced AI is inevitable and should be embraced. However, our framework suggests this accepts catastrophic risks that all actors would prefer to avoid if coordination were possible.
Break Condition 5 (Enforcement): Solve the enforcement problem through effective monitoring and verification mechanisms. Unlike software, which is inherently unverifiable, physical infrastructure (compute, energy, facilities) offers potential verification points. However, as discussed in the previous section, these advantages face structural erosion. Effective enforcement would require not merely technical verification capabilities but sustained international cooperation to prevent the competitive dynamics that undermine physical chokepoints, a coordination challenge as severe as the underlying AI governance problem itself.
The analysis suggests AI governance requires interventions of unprecedented scope and coordination. Breaking any single condition might suffice in principle, but the amplified intensity across all five conditions makes each intervention structurally more difficult than for historical arms control cases. This is not pessimism but structural realism: understanding the barriers is prerequisite to overcoming them.
@@ 5.6 Reconciliation with Conservative Productivity Estimates
Acemoglu (2025) provides conservative estimates of AI’s total productivity impact (approximately 0.55% annual growth over ten years). However, our framework operates independently of magnitude: even modest advantages create competitive pressure under competitive conditions because, mathematically speaking, the gap between two exponential growths widens with time, meaning economic and military advantages due to AI compound. Even conservative estimates show AI confers sufficient advantages that actors cannot afford restraint while competitors advance.
Uncertainty about AI’s potential also amplifies this: actors race based on potential upside or fear of loss rather than strict expected value. Consistent with Prospect Theory (Kahneman and Tversky, 1979), actors facing the ’sure loss’ of strategic obsolescence will gamble on risky acceleration rather than accept a certain disadvantage.
@@ 6.1 What the Framework Predicts
6.2 Structural Pressure vs. Metaphysical Necessity
This is not a deterministic claim about inevitable catastrophe. Rather, it identifies structural pressures creating strong tendencies toward particular outcomes. A ball will roll downhill unless an opposing force is applied.
What’s Required for Escape:
Successfully escaping the tragedy requires eliminating at least one of the five conditions:
Option 1—Reduce N: Consolidate actors to enable coordination. For productivity: reduce number of competing firms (faces antitrust concerns). For AI: consolidate development in single regulated entity (faces feasibility and risk concerns).
Option 2—Eliminate Externalities: Make individual choices not impose negative externalities on others. For both domains, this appears structurally impossible—competition inherently involves externalities.
Option 3—Break Dominance: Reduce existential stakes making defection dominant. For productivity: social safety nets, basic income, reduced competition, limit on capital mobility. For AI: reduce economic and military stakes (unclear how). This is perhaps most feasible.
Option 4—Accept Equilibrium: Give up on coordination, accepting mutual defection as outcome. For productivity: accept that leisure won’t increase. For AI: accept development risks as unavoidable.
Option 5—Solve Enforcement: Create effective monitoring and enforcement mechanisms. For productivity: global labor standards with verification and sanctions. For AI: international monitoring of compute and capabilities with enforcement. This is theoretically possible but practically difficult given international anarchy.
This is not pessimism but structural analysis. Understanding the structure is prerequisite for changing it. Interventions that ignore structural conditions will fail; interventions that address structural conditions have prospects for success.
@@ 6.3 Scope and Limitations
Temporal variation: The intensification of the tragedy after 1980 coincides with globalization, union decline, and increased capital mobility strengthening the conditions our framework identifies. However, detailed historical analysis is needed to establish definitive causal mechanisms for these temporal shifts.
Cross-national variation: Different institutional arrangements produce different manifestations of the productivity tragedy. Some countries exhibit the tragedy through stagnant wages despite productivity growth, others through stable hours but higher unemployment, and still others through pressure on benefits and job security. Our framework predicts that productivity gains will not flow fully to workers when competitive conditions are strong, but does not predict which specific dimension will bear the adjustment cost in each institutional context. Analysis of these institutional details and their interaction with competitive pressure is beyond our scope.
Condition intensity: As we have shown with the tragedy intensity extension, real-world conditions vary in intensity. When conditions hold weakly rather than strongly, partial coordination may be more feasible and sustainable. For instance, European labor market coordination succeeds in maintaining lower working hours than the United States, suggesting that weaker global competitive pressure (in non-tradable sectors) or stronger institutions can attenuate (though not eliminate) the tragedy. Future research should investigate systematically how condition intensity relates to the degree of coordination difficulty and the sustainability of partial coordination arrangements.
Empirical validation: We provide initial empirical evidence through declining labor share of income, productivity-wage decoupling, the European case for productivity, and the Russia-Ukraine drone war for AI development. However, comprehensive testing across historical periods, countries, sectors, and coordination attempts is necessary to fully validate our framework and identify boundary conditions. We view this paper as establishing a diagnostic framework and providing suggestive evidence, with systematic empirical validation remaining an important direction for future research.
@@ 7.3 Future Research
Identification of new tragedies: By checking the five conditions, can new coordination failures be identified in new domains?
Intensity measurement: What is the best way to quantify intensity? Is our extension a good starting point? What is the best formula to compute it and what does the best mean? How does intensity help in predictions of coordination difficulty?
Dynamic Modeling: How do these systems evolve? Does a tragedy in one domain (Productivity) accelerate the tragedy in another (AI) by providing the resources for the arms race?
Comparative governance analysis: Would a systematic comparison of governance attempts across domains reveal which structural conditions most resist coordination and which interventions most effectively target them?
Historical validation: Could an analysis of past coordination successes and failures through the framework reveal whether structural changes preceded successful coordination?

### v3

@@ 5.5 What Would Work?
Reduce winner-take-all dynamics and existential stakes making defection dominant. Mechanisms include: (1) credible international agreements reducing security competition between major powers, (2) economic restructuring through profit-sharing mechanisms such as the ‘Windfall Clause’ proposed by O’Keefe et al. (2020), where AI developers pre-commit to distributing exceptional profits (exceeding 1% of world GDP) broadly, and (3) demonstrated proof that advanced AI poses catastrophic risks even to prospective winners, undermining the perceived advantage of racing.
However, all three mechanisms face substantial barriers. International security agreements face severe commitment credibility problems: nations cannot reliably forego military advantages from transformative AI when adversaries might defect, creating overwhelming pressure to maintain competitive development. Economic restructuring through windfall clauses requires binding pre-commitment before competitive pressures emerge, making voluntary adoption difficult, and cannot address non-economic strategic advantages—military capability and geopolitical influence remain winner-take-all even with profit-sharing. Demonstrated proof of catastrophic risks faces an epistemic catch-22: definitive proof requires observing catastrophic outcomes, at which point it is too late. As Bostrom (2014) notes, the difficulty of credibly demonstrating novel existential risks before they materialize is a fundamental challenge, and even acknowledged risks may be insufficient when individual actors face dominant incentives to proceed if they believe others will do so regardless.
Break Condition 4 (Pareto-Inefficiency):
Accept the racing equilibrium as optimal, rejecting the premise that coordination would improve outcomes. Some argue that rapid AI deployment’s benefits outweigh coordination costs, effectively denying the Pareto-inefficiency condition.
However, this position faces severe challenges. As Bostrom (2014) and Russell (2019) document, the existential risks from misaligned superintelligence are non-excludable since even prospective winners cannot escape consequences of systems that exceed human control. The astronomical disvalue of existential catastrophe dominates even enormous competitive gains. This position requires either underestimating these non-excludable risks or assuming they are negligible, both of which the substantial AI safety research literature finds unconvincing. As discussed in Section 5.1, even prospective winners face negative expected utility from racing when probability of catastrophic outcomes is considered.
Break Condition 5 (Enforcement Difficulty):
Implement hardware-based verification using advanced semiconductor manufacturing as a chokepoint. Anderljung et al. (2023) argue that ‘compute governance’, e.g., monitoring and controlling access to high-end GPU clusters required for frontier AI training, offers the most viable enforcement mechanism, as specialized hardware is more detectable than software development. A global regime tracking chip production, data center construction, and computing cluster deployment could theoretically enforce development limits.
However, this advantage faces systematic erosion. As Sastry et al. (2024) document, algorithmic efficiency improvements systematically lower hardware requirements, allowing actors to achieve frontier capabilities with previously sub-frontier hardware—as demonstrated by DeepSeek’s early 2025 breakthrough achieving frontier results with dramatically reduced compute. Over multi-year timescales, hardware governance faces obsolescence as efficiency improvements enable powerful AI development on widely available hardware. Additionally, effective governance requires unprecedented international cooperation among competing powers, including adversarial nations with strong incentives to defect for strategic advantage. Verification of compliance faces challenges from concealed computing clusters and clandestine development programs.
Comprehensive Solution:
Resolving the AI governance tragedy requires interventions that simultaneously address multiple conditions: reducing the number of independent actors (C1), implementing credible verification mechanisms (C5), and reducing winner-take-all dynamics (C3). The most feasible approach likely combines:
  • International consortium among leading AI powers (partial C1 consolidation)
  • Compute governance with hardware tracking (partial C5 enforcement)
  • Windfall clauses and benefit-sharing mechanisms (partial C3 reduction)
  • Continued investment in technical AI safety (risk mitigation if governance fails)
However, our framework reveals the structural difficulty: each intervention faces severe implementation barriers, and partial solutions leave the tragedy substantially intact. The gap between theoretical solutions and practical feasibility measures the unprecedented nature of the coordination challenge. Unlike nuclear weapons, where physical constraints enabled partial success through non-proliferation regimes, AI’s software nature, dual-use characteristics, and low entry barriers create coordination requirements beyond historical precedent. The framework suggests we face a structurally harder problem than any successfully managed global coordination challenge to date.
@@ 5.6 Reconciliation with Conservative Productivity Estimates
Acemoglu (2025) provides conservative estimates of AI’s productivity impact (approximately 0.55% annual growth over ten years), attributing this limited potential to the prevalence of “so-so automation” that displaces labor without maximizing output. However, the structural tragedy operates independently of both magnitude and technological type. Even if AI offers only modest, displacement-heavy advantages, the dominance condition (C3) holds: firms cannot afford to forego even marginal efficiencies when competitors adopt them. The gap between exponential growth rates widens over time regardless of the base rate, meaning economic and military advantages compound. Conservative estimates still show AI confers sufficient advantages that actors cannot afford restraint while competitors advance.
Uncertainty amplifies this dynamic. Consistent with Prospect Theory (Kahneman and Tversky, 1979), actors facing the perceived certainty of strategic obsolescence will gamble on risky acceleration rather than accept certain disadvantage.
@@ 6.2 Scope and Limitations
Temporal Variation:
The intensification of the tragedy after 1980 coincides with globalization, union decline, and increased capital mobility strengthening the conditions our framework identifies. However, detailed historical analysis is needed to establish definitive causal mechanisms for these temporal shifts.
Cross-national Variation:
Different institutional arrangements produce different manifestations of the productivity tragedy. Some countries exhibit the tragedy through stagnant wages despite productivity growth, others through stable hours but higher unemployment, and still others through pressure on benefits and job security. Our framework predicts that productivity gains will not flow fully to workers when competitive conditions are strong, but does not predict which specific dimension will bear the adjustment cost in each institutional context. Analysis of these institutional details and their interaction with competitive pressure is beyond our scope.
Condition Intensity:
As we have shown with the tragedy intensity extension, real-world conditions vary in intensity. When conditions hold weakly rather than strongly, partial coordination may be more feasible and sustainable. For instance, European labor market coordination succeeds in maintaining lower working hours than the United States, suggesting that weaker global competitive pressure (in non-tradable sectors) or stronger institutions can attenuate (though not eliminate) the tragedy. Future research should investigate systematically how condition intensity relates to the degree of coordination difficulty and the sustainability of partial coordination arrangements.
Empirical Validation:
We provide initial empirical evidence through declining labor share of income, productivity-wage decoupling, the European case for productivity, and the Russia-Ukraine drone war for AI development. However, comprehensive testing across historical periods, countries, sectors, and coordination attempts is necessary to fully validate our framework and identify boundary conditions. We view this paper as establishing a diagnostic framework and providing suggestive evidence, with systematic empirical validation remaining an important direction for future research.
@@ 7.3 Future Research
Identification of New Tragedies:
By checking the five conditions, can new coordination failures be identified in emerging domains? Do technologies like quantum computing, synthetic biology, or geoengineering exhibit the structural conditions that predict coordination difficulty? Could applying the framework to existing problems where solutions have failed reveal structural barriers previously overlooked, suggesting interventions that target the underlying conditions rather than symptoms?
Intensity Measurement:
What is the best way to quantify condition intensity? Is our Tragedy Index formulation a useful starting point, or would alternative aggregation methods better capture coordination difficulty? How does measured intensity correlate with observed governance outcomes across historical cases? Could systematic application of game-theoretic measures (e.g., Price of Anarchy) across domains validate the Index’s ordinal rankings and reveal whether structural and payoff-based approaches converge?
Formalization Refinement:
Our framework applies game-theoretic concepts largely heuristically. Could dynamic models with continuous strategies and empirically calibrated parameters confirm the tragedy persists under more realistic assumptions? Would computational simulations of multi-agent systems validate equilibrium predictions?
Productivity Model Extensions:
Our productivity formalization uses simplified utility functions and assumes comparable productivity gains. Could more realistic utility specifications incorporating risk aversion, time preferences, and worker-firm bargaining better capture the mechanisms? How does the tragedy structure change under persistent heterogeneous productivity shocks where market consolidation is incomplete or ongoing?
Cross-domain Dynamics:
How do tragedies in different domains interact? Does the productivity tragedy accelerate AI development by providing resources and competitive pressure for the capabilities race? Can success in coordinating one domain create spillovers that facilitate coordination in structurally similar domains?
Comparative Governance Analysis:
Would systematic comparison of governance attempts across domains reveal which structural conditions most resist coordination? Which interventions most effectively target specific conditions, and under what circumstances do partial solutions succeed or fail?
Historical Validation:
Could analysis of past coordination successes and failures through the framework reveal whether structural changes preceded successful coordination? Do cases like the Montreal Protocol, nuclear test bans, or labor regulations show condition-breaking interventions, or merely temporary overrides that eventually eroded?
Military AI and Conflict Escalation:
Does autonomous weapons development create a distinct tragedy where AI-enabled military systems lower the threshold for initiating conflict? If decision-making speed and reduced human casualties make war less costly for attackers, does this break the mutual restraint that deterrence theory assumes? Could the framework predict which military applications of AI face the most severe coordination barriers?
Institutional Resilience:
The European case shows partial coordination under favorable conditions, but with signs of erosion. What determines the sustainability of costly institutional overrides? Can we predict which coordination arrangements will persist versus collapse under competitive pressure?
Implications for Political Science and Economics:
Could the structural tragedy framework unify phenomena currently treated separately across disciplines? Would reframing problems like arms races, trade wars, regulatory competition, or financial contagion through the five-condition lens reveal common structural solutions? Can the framework inform institutional design in international relations and political economy?